# Locally integrable cross sections and their intersection covolume

### v2

@@ 1.1. Cross sections, intensity and transverse measures
Cross sections provide a way to reduce the continuous dynamics of a locally compact second countable (lcsc) unimodular group to a countable equivalence relation, while retaining key quantitative invariants. They appear naturally in point processes, ergodic theory, aperiodic order, and other related fields. Formally, let $G$ be an lcsc unimodular group acting in a measure preserving way on a standard probability space $\left(X,\mu\right)$. A cross section is a Borel set $Y\subset X$ meeting every $G$-orbit, such that each return times set
@@
and $Y$ is locally integrable if $\iota_{\mu}\left(Y\right)<+\infty$, in which case we speak of a transverse $G$-space $(X,\mu,Y)$.
@@
The construction of transverse measures for flows (actions of $\mathbb{R}$) dates back to Ambrose–Kakutani (see [21, §12]). For unimodular groups, separated cross sections (those with $\Lambda_{Y}$ discrete) have been studied in [3, 16, 22, 7] among others. Here we develop a unified construction for arbitrary cross sections, adapting Palm theory to the ergodic setting and extending the intersection space framework due to the second author, Hartnick and Karasik [7].
A key step is showing that the classical Mecke equation characterizes the right notion of invariance for transverse measures. This confirms a conjecture of G. Last [17, §3.10], and leads to a canonical transverse correspondence:
@@
This correspondence will be central to what follows.
@@ 1.2. Periodicity and intersection covolume
We now consider the following (minimal) notion of periodicity for cross sections: a cross section $Y$ is called completely periodic if each return times set $Y_{x}$ is a coset of a fixed lattice $\Gamma<G$. In this case, the action factors through the homogeneous space $\left(\Gamma\backslash G,m_{\Gamma\backslash G}\right)$, and $Y$ collapses to a single point in the factor. More generally, for a transverse $G$-space $\left(X,\mu,Y\right)$, we seek to quantify the deviation of $Y$ from complete periodicity.
To this end, we use the intersection space construction presented in [7], adapted to general cross sections using the aforementioned transverse correspondence. We then define, for a transverse $G$-space $\left(X,\mu,Y\right)$, the intersection covolume,
$I_{\mu}\left(Y\right)\in\left(0,+\infty\right],$
to be the total mass of the intersection measures (see Section 7).
@@
In Section 5.2, we characterize transverse $G$-factors also in terms of transverse measures. Importantly, we show that both the intensity and the intersection covolume are preserved under transverse $G$-factors (see Propositions 5.10 and 7.5).
@@
A probability preserving $G$-space $\left(X,\mu\right)$ admits a cross section $Y$ with $I_{\mu}\left(Y\right)=\iota_{\mu}\left(Y\right)$ if and only if it is an induced action from a lattice in $G$.
The intersection covolume $I_{\mu}\left(Y\right)$ can be infinite even when $\iota_{\mu}\left(Y\right)<+\infty$. For instance, the Poisson point process on $\mathbb{R}$ has finite intensity, while its intersection covolume can be shown to be infinite. It is therefore of particular interest to study situations where the inequality in Theorem A is strict, yet the intersection covolume remains finite, thus indicating a degree of periodicity without being completely periodic. With a simple use of suspension flows (flows under a function), we will show in Section 8.1 that for every $\epsilon>0$, there is an ergodic transverse $\mathbb{R}$-space $\left(X,\mu,Y\right)$ such that
$\iota_{\mu}\left(Y\right)<I_{\mu}\left(Y\right)<\left(1+\epsilon\right)\cdot\iota_{\mu}\left(Y\right).$
@@ 1.3. Cut–and–project actions and their cross sections
Cut–and–project schemes arise in the study of quasicrystals and aperiodic order, where cut–and–project sets provide ordered but non-periodic point patterns in groups. For uniform cut–and–project schemes (i.e., when the underlying lattice is uniform), it was proved in [7, Theorem 1.4] that the intersection measure is finite. In the following, we extend this result to nonuniform lattices, and thus their intersection spaces constitute new examples of cross sections with finite intersection covolume but which are far from completely periodic.
@@
as a Borel $G$-space ($G$ acts on the first coordinate). For a window $W\subset H$, define the cross section
@@ 1.4. Groups without lattices
Let $G=\mathbb{Q}_{p}$ or $G=\mathbb{A}_{\mathrm{fin}}$. For every probability preserving $G$-space $\left(X,\mu\right)$ with a cross section $Y$ such that $\Lambda_{Y}$ is uniformly discrete, it holds that
@@
A family of groups without lattices introduced in [6, §2.4] is given by the one-parameter family of nilpotent Lie groups $G_{\lambda}$, $\lambda\in\left(0,+\infty\right)$. It was shown there that $G_{\lambda}$ admits a lattice if and only if $\lambda$ is rational, and that $G_{\lambda}$ admits cut–and–project actions when $\lambda$ is an algebraic irrational. It was shown in [18, Theorem 1.5] that $G_{\lambda}$ admits no cut–and–project action when $\lambda$ is transcendental. This motivates the following problem:
@@ 2. General preliminaries
For a standard Borel space $X$, we denote by $\mathcal{M}\left(X\right)$ be the set of Borel $\sigma$-finite measures on $X$. For $\mu\in\mathcal{M}\left(X\right)$ and a Borel function $f:X\to\left[0,+\infty\right]$ we write
@@ 3.2. Basic definitions
Let us start by defining measure invariance to countable Borel equivalence relations. For the basics of this notion of measure invariance see [15, §8], [14, §4.2 & 4.7]. Let $A$ be a standard Borel space and $E$ a countable Borel equivalence relation on $A$. Thus, $A$ is a Borel subset of $A\times A$, the relation $\left(a,a^{\prime}\right)\in E$ for $a,a^{\prime}\in A$ is an equivalence relation, and each $E$-class $E\left(a\right)\coloneqq\{a^{\prime}\in A:\left(a,a^{\prime}\right)\in E\}$ is at most countable. We associate with $E$ the set (pseudo-group) $\bigl[\!\bigl[E\bigr]\!\bigr]$ of all partial transformations
@@
Thus, $E_{G}^{Y}$ is the restriction to $Y$ of the orbit equivalence relation induced by the action of $G$ on $X$. Since $Y$ is a cross section, $E_{G}^{Y}$ is indeed countable, i.e. each $E_{G}^{Y}$-class is countable.
@@
Our proof of Theorem 3.4 uses crucially the Lusin–Novikov uniformization theorem. We use the following form of this classical theorem that appears in [13, Theorem (18.10)]. [^2] Note that by the first part of [13, Theorem (18.10)], every Borel graph is an actual graph of a Borel map. Then the following is a reformulation of the Moreover part of [13, Theorem (18.10)].
@@
The first way we rely on the Lusin–Novikov theorem 3.5 is the following well-known reformulation of the general notion of invariance of measures to countable Borel equivalence relations. For a proof (relying on the Lusin–Novikov theorem) see [15, §8].
@@
and hence is countable and nonempty. By the Lusin–Novikov theorem 3.5, [^3] Here we only use the first part of [13, Theorem (18.10)]. there is a Borel map $\lambda:\mathrm{Dom}\left(\tau\right)\to G$ such that $\tau\left(y\right)=\lambda\left(y\right).y$ for all $y\in\mathrm{Dom}\left(\tau\right)$. Define then the Borel function
@@ 4. The transverse correspondence
We can now formulate the transverse correspondence, and to this end we need first to identify the range of the transverse measure construction, namely what measures arise as transverse measures. In ergodic-theoretical context, the natural notion of invariance is the general notion of measures being invariant to countable Borel equivalence relations as defined above.
@@
Note that the transverse correspondence 4.3 is subject to a specific choice of Haar measure on $G$.
@@ 4.1. Proof of Campbell theorem
We now deduce that the range of the transverse correspondence 4.3 is indeed $\mathcal{M}^{E_{G}^{Y}}\left(Y\right)$. To this end we argue similarly to along the lines of [17, Theorem 3.15] and exploit Theorem 3.4.
@@ 4.2. Injective covers
Let $X^{\prime}$ be a Borel $G$-space with a cross section $Y^{\prime}$, and let $X\subseteq X^{\prime}$ be a Borel $G$-subspace (i.e. a $G$-invariant Borel subset) with a cross section $Y$, such that $Y\subseteq Y^{\prime}$. Then there are Borel partitions of unity, $\rho:G\times X\to\left[0,1\right]$ for $X$ with respect to $Y$ and $\rho^{\prime}:G\times X^{\prime}\to\left[0,1\right]$ for $X^{\prime}$ with respect to $Y^{\prime}$, such that
@@
is measure preserving, i.e. it pushes $\left(m_{G}\otimes\mu_{Y}\right)\mid_{C}$ to $\mu\mid_{\mathrm{a}\left(C\right)}$.
@@ 4.3. Final proof of the transverse correspondence
In light of Proposition 4.4 and Theorem 3.4, we need to show that the correspondence $\mathcal{M}^{G}\left(Y\right)\to\mathcal{M}_{\mathrm{Mecke}}^{G}\left(Y\right)$, $\mu\mapsto\mu_{Y}$, forms a bijection. The injectivity has been established in Proposition 4.10. For the surjectivity, for every $\nu\in\mathcal{M}_{\mathrm{Mecke}}^{G}\left(Y\right)$ we have constructed $\nu^{X}\in\mathcal{M}^{G}\left(X\right)$ via the correspondence (4.3.1). By Claim 4.11, $\nu^{X}\left(f_{X}\right)=m_{G}\otimes\nu\left(f\right)$ for every Borel function $f:G\times Y\to\left[0,+\infty\right]$, and by Proposition 4.10, this property determines $\nu_{Y}^{X}$ uniquely among the elements of $\mathcal{M}_{\mathrm{Mecke}}^{G}\left(Y\right)$, which include $\nu$, concluding that $\nu_{Y}^{X}=\nu$. ∎
@@ 5. Transverse $G$-spaces and $G$-factors
$\Lambda_{Y}\coloneqq\{g\in G:g.Y\cap Y\}=\bigcup\nolimits_{y\in Y}Y_{y}.$
Note that $\Lambda_{Y}$ is symmetric (i.e. $\Lambda_{Y}^{-1}=\Lambda_{Y}$), and we will refer to $\Lambda_{Y}$ as the return times set of $Y$. Then a cross section $Y$ of a Borel $G$-space $X$ is said to be separated, if $e_{G}\in\Lambda_{Y}$ is an isolated point. The proof of the following lemma is elementary (see [7, Lemma 3.1(ii)]).
@@ 5.2. Transverse $G$-factors
establishing (5.2.1). Using [26, Lemma B.8], [^7] This lemma follows from Varadarajan’s compact model theorem [24, Ch. V, §3, Theorem 5.7], with a theorem attributed to Kallman [26, Theorem A.5], which is the uniformization theorem [13, Theorem (18.18)]. there is a Borel set $X_{o}^{o}\subseteq X^{o}$ such that $\mu\left(X^{o}\setminus X_{o}^{o}\right)=0$ and $G.X_{o}^{o}$ is Borel, as well as a Borel map $\lambda:G.X_{o}^{o}\to G$ such that $\lambda\left(x\right).x\in X_{o}^{o}$ for every $x\in G.X_{o}^{o}$ and $\lambda\left(x\right)=e_{G}$ for every $x\in X_{o}^{o}$. Put the Borel set

### v3

@@ 1. Introduction
This paper, together with its companion paper [4], develops a systematic theory of intersection spaces associated with cross sections of probability preserving actions of locally compact second countable (lcsc) unimodular groups. Intersection spaces were originally introduced by the second author together with Hartnick and Karasik in [7], where their construction relied on the additional assumption that the cross sections are separated. The present work removes this restriction and treats general locally integrable cross sections, thereby significantly enlarging the range of applications. In particular, we can treat random subsets of general, not necessarily uniform, approximate lattices (in particular those arising from cut–and–project schemes), extending the scope of the theory well beyond the framework accessible in [7]. In another direction, the framework developed in this work has been used by Hartnick and Sarti [10] to establish bounded cohomological induction for transverse measured groupoids.
By considering general locally integrable cross sections, we unify the transverse measure construction for separated cross sections in ergodic theory with the Palm measure construction in point process theory. Accordingly, a key ingredient in our framework is a characterization of relative invariance in terms of the classical Mecke equation. This was known for lcsc abelian groups [11, 12], and by proving it for general lcsc groups we confirm a conjecture of G. Last [17, §3.10], which is of independent interest already in point processes theory.
Intersection spaces were tailored to the study of random subsets of approximate lattices in the sense of [6]. Although the construction extends to arbitrary locally integrable cross sections, the canonically associated intersection space measure is typically infinite. Finiteness of the intersection space measure arises only in structurally constrained situations, most prominently in the presence of approximate lattice phenomena or after a suitable thinning. In these cases, its total mass, which we call the intersection covolume, serves as a quantitative measure of periodicity of the system. This interpretation is made precise in Theorem A below, whose proof is based on a new formula for the intersection covolume as a higher order version of Kac’s lemma (see Section 7).
@@ 1.1. Cross sections, intensity and transverse measures
Cross sections provide a way to reduce the continuous dynamics of an lcsc unimodular group to a countable equivalence relation, while retaining key quantitative invariants. They appear naturally in point processes, ergodic theory, aperiodic order, and other related fields. Formally, let $G$ be an lcsc unimodular group acting in a measure preserving way on a standard probability space $\left(X,\mu\right)$. A cross section is a Borel set $Y\subset X$ meeting every $G$-orbit, such that each return times set
@@
and $Y$ is locally integrable if $\iota_{\mu}\left(Y\right)<+\infty$, in which case we speak of a transverse $G$-space $\left(X,\mu,Y\right)$.
@@
The construction of transverse measures for flows (actions of $\mathbb{R}$) dates back to Ambrose–Kakutani (see [21, §12]). For unimodular groups, separated cross sections (those with $\Lambda_{Y}$ discrete) have been studied in [3, 16, 22, 7] among others. In order to extend the intersection spaces construction due to [7] to general cross sections, we need to extend the transverse measure construction accordingly. To this end we prove that the Mecke equation characterizes relative invariance of transverse measures, thereby confirming Last’s conjecture [17, §3.10], and this leads to a canonical transverse correspondence:
@@ 1.2. Periodicity and intersection covolume
We now consider the following (minimal) notion of periodicity for cross sections: a cross section $Y$ is called completely periodic if each return times set $Y_{x}$ is a coset of a fixed lattice $\Gamma<G$. In this case, the action factors through the homogeneous space $\left(\Gamma\backslash G,m_{\Gamma\backslash G}\right)$, and $Y$ collapses to a single point in the factor. More generally, for a transverse $G$-space $\left(X,\mu,Y\right)$, we seek to quantify the deviation of $Y$ from complete periodicity. To this end, we use the intersection space construction presented in [7], adapted to general cross sections using the aforementioned transverse correspondence.
Let then $\left(X,\mu\right)$ be probability preserving $G$-space with a locally integrable cross section $Y$, and let $\mu_{Y}$ be its transverse measure. The associated intersection space (of order $2$) is the measure preserving $G$-space $\left(Y^{\left[2\right]},\mu^{\left[2\right]}\right)$ defined as follows. The underlying Borel $G$-space is
$Y^{\left[2\right]}:=\{\left(g.y_{1},g.y_{2}\right):g\in G,y_{1},y_{2}\in Y\},$
with the diagonal action of $G$. One notes that $Y\times Y$ serves as a cross section of $Y^{\left[2\right]}$, and therefore, using the transverse correspondence, the $\sigma$-finite measure $\mu^{\left[2\right]}$ is the unique measure whose transverse measure with respect to $Y\times Y$ is the product transverse measure $\mu_{Y}\otimes\mu_{Y}$. We then define the intersection covolume (of order $2$) of $\left(X,\mu,Y\right)$ to be the quantity
$I_{\mu}\left(Y\right):=\mu^{\left[2\right]}\big(Y^{\left[2\right]}\big)\in\left(0,+\infty\right].$
This will be presented in details in Section 7, where we define intersection covolume $I_{\mu}^{r}\left(Y\right)$ for all $r\in\mathbb{N}$.
@@
In Section 5.2, we characterize transverse $G$-factors in terms of transverse measures. Importantly, both intensity and intersection covolume are preserved under transverse $G$-factors (Propositions 5.10 and 7.5).
@@
An ergodic probability preserving $G$-space $\left(X,\mu\right)$ is an induced action from a lattice in $G$ if and only if it admits a cross section $Y$ satisfying $I_{\mu}\left(Y\right)=\iota_{\mu}\left(Y\right)$.
The intersection covolume $I_{\mu}\left(Y\right)$ can be infinite even when $\iota_{\mu}\left(Y\right)<+\infty$. For instance, the Poisson point process on $\mathbb{R}$ has finite intensity, while its intersection covolume can be shown to be infinite. It is therefore of particular interest to study situations where the inequality in Theorem A is strict, yet the intersection covolume remains finite, thus indicating a degree of periodicity without being completely periodic. With a simple use of suspension flows, we will show in Section 8.1 that for every $\epsilon>0$, there is an ergodic transverse $\mathbb{R}$-space $\left(X,\mu,Y\right)$ such that $\iota_{\mu}\left(Y\right)<I_{\mu}\left(Y\right)<\left(1+\epsilon\right)\cdot\iota_{\mu}\left(Y\right)$.
@@ 1.3. Cut–and–project actions and their cross sections
Cut–and–project schemes arise in the study of quasicrystals and aperiodic order, where cut–and–project sets provide ordered but non-periodic point patterns in groups. For uniform cut–and–project schemes (when the underlying lattice is uniform), it was proved in [7, Theorem 1.4] that the intersection measure is finite. We extend this to nonuniform lattices, and thus their intersection spaces constitute new examples of cross sections with finite intersection covolume but which are far from completely periodic.
@@
as a Borel $G$-space ($G$ acts on the first coordinate). A measurable set $W\subset H$ is called a window, if it is relatively compact, has nonempty interior, and it is aperiodic in the sense that no nontrivial element of $H$ stabilizes $W$. For a window $W\subset H$, define a cross section of $\Xi$ by
@@ 1.4. Groups without lattices
Let $G=\mathbb{Q}_{p}$ or $G=\mathbb{A}_{\mathrm{fin}}$. For every ergodic probability preserving $G$-space $\left(X,\mu\right)$ with a cross section $Y$ such that $\Lambda_{Y}$ is uniformly discrete, it holds that
@@
A class of groups without lattices introduced in [6, §2.4] is the one-parameter family of nilpotent Lie groups $G_{\lambda}$, $\lambda\in\left(0,+\infty\right)$. It was shown there that $G_{\lambda}$ admits a lattice if and only if $\lambda$ is rational, and that $G_{\lambda}$ admits cut–and–project actions when $\lambda$ is an algebraic irrational. It was shown in [18, Theorem 1.5] that $G_{\lambda}$ admits no cut–and–project action when $\lambda$ is transcendental. This motivates the following:
@@ 2. General preliminaries
For a standard Borel space $X$, we denote by $\mathcal{M}\left(X\right)$ the set of Borel $\sigma$-finite measures on $X$. For $\mu\in\mathcal{M}\left(X\right)$ and a Borel function $f:X\to\left[0,+\infty\right]$ we write
@@ 3.2. Basic definitions
Let us start by defining measure invariance to countable Borel equivalence relations. For the basics of this notion of measure invariance see [13, §8], [15, §4.2 & 4.7]. Let $A$ be a standard Borel space and $E$ a countable Borel equivalence relation on $A$. Thus, $E$ is a Borel subset of $A\times A$, the relation $\left(a,a^{\prime}\right)\in E$ for $a,a^{\prime}\in A$ is an equivalence relation, and each $E$-class $E\left(a\right)\coloneqq\{a^{\prime}\in A:\left(a,a^{\prime}\right)\in E\}$ is at most countable. We associate with $E$ the set (pseudo-group) $\bigl[\!\bigl[E\bigr]\!\bigr]$ of all partial transformations
@@
Thus, $E_{G}^{Y}$ is the restriction to $Y$ of the orbit equivalence relation induced by the action of $G$ on $X$. Since $Y$ is a cross section, $E_{G}^{Y}$ is indeed countable, namely each $E_{G}^{Y}$-class is countable.
@@
Our proof of Theorem 3.4 uses crucially the Lusin–Novikov uniformization theorem. We use the following form of this classical theorem that appears in [14, Theorem (18.10)]. [^2] Note that by the first part of [14, Theorem (18.10)], every Borel graph is an actual graph of a Borel map. Then the following is a reformulation of the Moreover part of [14, Theorem (18.10)].
@@
The first way we rely on the Lusin–Novikov theorem 3.5 is the following well-known reformulation of the general notion of invariance of measures to countable Borel equivalence relations. For a proof (relying on the Lusin–Novikov theorem) see [13, §8].
@@
and hence is countable and nonempty. By the Lusin–Novikov theorem 3.5, [^3] Here we only use the first part of [14, Theorem (18.10)]. there is a Borel map $\lambda:\mathrm{Dom}\left(\tau\right)\to G$ such that $\tau\left(y\right)=\lambda\left(y\right).y$ for all $y\in\mathrm{Dom}\left(\tau\right)$. Define then the Borel function
@@ 4. The transverse correspondence
We now formulate the transverse correspondence, and to this end we recall the set $\mathcal{M}^{E_{G}^{Y}}\left(Y\right)$ of all $E_{G}^{Y}$-invariant Borel $\sigma$-finite measures on $Y$.
@@
Note that the transverse correspondence depends on a specific normalization of $G$’s Haar measure.
@@ 4.1. Proof of Campbell theorem
We now deduce that the range of the transverse correspondence 4.3 is indeed $\mathcal{M}^{E_{G}^{Y}}\left(Y\right)$. To this end we argue similarly to [17, Theorem 3.15] and exploit Theorem 3.4.
@@ 4.2. Injective covers
Let $X^{\prime}$ be a Borel $G$-space with a cross section $Y^{\prime}$, and let $X\subseteq X^{\prime}$ be a Borel $G$-subspace (i.e., a $G$-invariant Borel subset) with a cross section $Y$, such that $Y\subseteq Y^{\prime}$. Then there are Borel partitions of unity, $\rho:G\times X\to\left[0,1\right]$ for $X$ with respect to $Y$ and $\rho^{\prime}:G\times X^{\prime}\to\left[0,1\right]$ for $X^{\prime}$ with respect to $Y^{\prime}$, such that
@@
is measure preserving, meaning that it pushes $\left(m_{G}\otimes\mu_{Y}\right)\mid_{C}$ to $\mu\mid_{\mathrm{a}\left(C\right)}$.
@@ 4.3. Final proof of the transverse correspondence
In light of Proposition 4.4 and Theorem 3.4, we need to show that the correspondence $\mathcal{M}^{G}\left(X\right)\to\mathcal{M}_{\mathrm{Mecke}}^{G}\left(Y\right)$, $\mu\mapsto\mu_{Y}$, forms a bijection. The injectivity has been established in Proposition 4.10. For the surjectivity, for every $\nu\in\mathcal{M}_{\mathrm{Mecke}}^{G}\left(Y\right)$ we have constructed $\nu^{X}\in\mathcal{M}^{G}\left(X\right)$ via the correspondence (4.3.1). By Claim 4.11, $\nu^{X}\left(f_{X}\right)=m_{G}\otimes\nu\left(f\right)$ for every Borel function $f:G\times Y\to\left[0,+\infty\right]$, and by Proposition 4.10, this property determines $\nu_{Y}^{X}$ uniquely among the elements of $\mathcal{M}_{\mathrm{Mecke}}^{G}\left(Y\right)$, which include $\nu$, concluding that $\nu_{Y}^{X}=\nu$. ∎
@@ 5. Transverse $G$-spaces and $G$-factors
$\Lambda_{Y}\coloneqq\{g\in G:g.Y\cap Y\neq\emptyset\}=\bigcup\nolimits_{y\in Y}Y_{y}.$
Note that $\Lambda_{Y}$ is symmetric (namely $\Lambda_{Y}^{-1}=\Lambda_{Y}$), and we will refer to $\Lambda_{Y}$ as the return times set of $Y$. Then a cross section $Y$ of a Borel $G$-space $X$ is said to be separated, if $e_{G}\in\Lambda_{Y}$ is an isolated point. The proof of the following lemma is elementary (see [7, Lemma 3.1(ii)]).
@@ 5.2. Transverse $G$-factors
establishing (5.2.1). Using [26, Lemma B.8], [^7] This lemma follows from Varadarajan’s compact model theorem [24, Ch. V, §3, Theorem 5.7], with a theorem attributed to Kallman [26, Theorem A.5], which is the uniformization theorem [14, Theorem (18.18)]. there is a Borel set $X_{o}^{o}\subseteq X^{o}$ such that $\mu\left(X^{o}\setminus X_{o}^{o}\right)=0$ and $G.X_{o}^{o}$ is Borel, as well as a Borel map $\lambda:G.X_{o}^{o}\to G$ such that $\lambda\left(x\right).x\in X_{o}^{o}$ for every $x\in G.X_{o}^{o}$ and $\lambda\left(x\right)=e_{G}$ for every $x\in X_{o}^{o}$. Put the Borel set